25. Today (T-O), an investor purchased a 20 year bond with a 5.00% coupon and a face value of $100,000 for $106,550. In six months (T-0.5) interest rates have decreased by 0.50% and the investor decides to sell the bond immediately after receiving the first coupon payment. What is the investors total gain (loss) on the bond? HINT: Total Gain (Loss) Price Change in Bond +Coupon A. ($6,548) B. ($6,048) C. $7,130 D. $7,602 E. $7,630 Assume all future cash flows (FCFF, FCFE, and Dividends) will grow at a constant sustainable growth rate policy. The companys cost of debt is 20% while the companys cost of equity is 30%; the company has an effective tax rate of 35%. Use the following information in the table below to help answer problems 26-27: FCFE Today (T O) FCFF Today (T-0) Shareholder Equity Total Debt Total Assets Net Income Dividends Shares Outstanding in millions 800 750 400 600 1000 100 40 25 26. An activist investor who wants to purchase all the companys shares would be willing to pay approximately_ Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a. $165 per share b. $176 per share c. $213 per share d. $230 per share e. $245 per share 27. A private equity group who wants to purchase all of the companys assets would be willing to pay approximately . Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a. $9,583M b. $10,222M c. $10,518M d. $17,969M e. $19,167M
Answer #1

Question No. 25 is not answered.

Answers to questions 26 and 27:                                                   (All figures in millions)

Given, Net income(PAT) = 100        Dividend = 40                  No. of shares = 25      Share holder fund (SHF) = 400

Debt = 600               Cost of equity (Ke) = 30% Cost of debt (Kd) = 20% (1-0.35) = 13%   

Note: For calculating cost of debt, rate of interest net of taxes to be considered as there is savings to the extent of tax benefit available on interest expenses. Hence, the cost of debt of 13% and not 20%.

(i) Earning per share (EPS) = PAT / No. of shares

= 100/25

=4

(ii) Dividend per share(DPS) = Dividends / No. of shares

= 40/25

= 1.6

(iii) Retention ratio (b) = 1 - (DPS / EPS)

= 1 - ( 1.6 / 4 )

= 0.6 i.e., 60%

(iv) Return on Equity (r) = Net inocme / SHF

= 100 / 400

= 25%

(v) It is given in the question that growth rate (g) is dependent on dividend payout policy and return on equity. Hence, the formula to calculate growth rate is   g = b r

= 60% * 25%

= 15% . Therefore, the growth rate is 15%.

(vi) cost of capital of the company (k) =( Kd * Wd ) + (Ke * We)

Where Kd = cost of debt, Wd = Debt weightage, Ke = Cost of equity, We = Equity weightage.

Kd = 13% (calculated above),     Ke = 30% (given),    Wd = (Debt / Total fund) = 600 / 1000 = 60%

We = (Equity / Total fund) = 400/1000 = 40%

K = (Kd * Wd) + (Ke * We)

= (0.13 * 0.6) + (0.4 * 0.3)

= 0.078 + 0.12

= 0.198 i.e., 19.8%

Answer to Question 26:

This has been solved using FCFE as given in the question.

Formula to calculate value of the firm using FCFE approach is = FCFE (T=1) / ( Ke - g)

Where, FCFE (T=1) is FCFE after one year from now. which can be calculated as follow:

FCFE (1) = FCFE (T=0) + g

= 800 + 15%

= 920

Value of the company = FCFE (T=1) / (Ke - g)

= 920 / (0.3 - 0.15)

= 920 / 0.15

= 6,133.33 millions

Value per share = 6,133.33 / 25

= 245.33 i.e., 245

Hence, the investor who wants to purchase the shares of the company would pay $ 245 which is the value per share.

Answer to Question 27:

The answer is based on FCFF approach as given in the question.

Value of the company = FCFF (T=1) / (K - g)

Where FCFF (T=1) is FCFF of the company after one year from now. Which is calculated as below:

FCFF (T=1) = FCFF (T=0) + g

= 750 + 15%

= 862.5

Value of the company = FCFF (T=1) / (K - g)

= 862.5 / (0.198 - 0.15)

= 862.5 / 0.048

= 17968.75 i.e., 17,969 millions

Hence, the company which wants to buy the company's assets would be willing to pay $ 17,969 millions which is the value of the company.